OpenStudy (swagster23780):

Steven recorded the growth of a plant over 10 weeks for his science project. He made a graph that shows how much the plant grew each week. The plant was 2 in. tall when Steven started his project. At week 5, the plant was 8 in. tall. What is the equation that represents Steven’s situation?

OpenStudy (swagster23780):

A. y=5/6x+2 B. y=8/5x+2 C. y=6/5x+2 D. y=5/8x+2

OpenStudy (swagster23780):

Please Help

OpenStudy (mww):

to find the equation calculate the slope of the pts. You have two pts being (time, height) as A(0,2) and B(5,8) The slope (gradient) is given by \[\frac{ y2 - y1 }{ x2 - x1 }\] Plug the two pts in to find the slope. Afterwards pick ONE of the points, say A and then sub it into this equation along with the gradient you just found \[y - y_1 = m(x-x_1)\]

OpenStudy (mww):

as the only difference between the answers is the slope you actually only really need the slope to answer this MCQ.